# Can language models boost the power of randomized experiments without statistical bias?

### v1

@@ Abstract
Randomized experiments or randomized controlled trials (RCTs) are gold standards for causal inference, yet cost and sample-size constraints limit power. Meanwhile, modern RCTs routinely collect rich, unstructured data that are highly prognostic of outcomes but rarely used in causal analyses. We introduce CALM (Causal Analysis leveraging Language Models), a statistical framework that integrates large language models (LLMs) predictions with established causal estimators to increase precision while preserving statistical validity. CALM treats LLM outputs as auxiliary prognostic information and corrects their potential bias via a heterogeneous calibration step that residualizes and optimally reweights predictions. We prove that CALM remains consistent even when LLM predictions are biased and achieves efficiency gains over augmented inverse probability weighting estimators for various causal effects. In particular, CALM develops a few-shot variant that aggregates predictions across randomly sampled demonstration sets. The resulting U-statistic-like predictor restores i.i.d. structure and also mitigates prompt-selection variability. Empirically, in simulations calibrated to a mobile-app depression RCT, CALM delivers lower variance relative to other benchmarking methods, is effective in zero- and few-shot settings, and remains stable across prompt designs. By principled use of LLMs to harness unstructured data and external knowledge learned during pretraining, CALM provides a practical path to more precise causal analyses in RCTs.
Keywords: Artificial intelligence; Clinical trials; Few-shot learning; Unstructured data.
@@ 1.1 Background and motivation
Results obtained from reliably designed randomized experiments or randomized controlled trials (RCTs) are often considered to be evidence of the highest grade for assessing the effectiveness of biomedical or behavioral interventions. Through randomization and controlled conditions, RCTs minimize bias and allow reliable estimation of causal effects. As secondary analyses, subgroup analyses, the study of treatment effect heterogeneity, are also frequently conducted within RCTs to explore heterogeneity of causal effects and to support the development of precision medicine. However, RCTs are often constrained by limited sample sizes and insufficient power due to high implementation costs, particularly for detecting effects within subpopulations.
Furthermore, a growing number of RCTs across various fields now collect unstructured data to capture the rich contextual information often missed by structured variables, and these unstructured data can be highly predictive of experimental outcomes. Unstructured data, such as clinical notes, subject narratives, and transcripts of subject-provider interactions, are becoming standard in RCTs spanning oncology, mental health [45, 49], dementia [22], and healthcare delivery [11, 24] (see Table 1 in the Supplementary Materials for some motivating RCTs). As another example from our motivating case study, the BRIGHTEN study (an app-based therapy for depression as described in Section 5.1) illustrates a similar observation. Figure 1(A) summarizes the trial’s collected data, where we observe that unstructured data are collected at the initiation of the study. Figure 1(B) shows that baseline free-text responses on motivation for app use and follow-up app-satisfaction comments exhibit the highest cosine similarity with discretized PHQ-9 outcomes (trial primary outcome), surpassing demographics and app-usage metrics. Together, these findings demonstrate the potential of unstructured data as valuable “prognostic factors” (pre-treatment variables that are predictive of the outcome), for which existing literature documents that adjusting for such variables can substantially enhance the precision and power of RCT analyses [8, 29, 62, 63]. Leveraging unstructured data, therefore, represents an opportunity to enhance causal estimation and detect treatment heterogeneity in RCTs.
@@
Recent advances in artificial intelligence (AI), large-scale pre-trained large language models (LLMs) across different data modalities, in particular, offer potential opportunities to boost the statistical power of RCTs by leveraging the external knowledge obtained during training and by extracting additional information from unstructured data. LLMs are deep neural networks trained with self-supervision on broad and heterogeneous corpora, producing general representations that can be rapidly adapted to many downstream tasks with little or no additional labeled data [9]. Multimodal LLMs now span the biomedical and behavioral spectrum, with representative examples including GPT-4 [42], Gemini [57], and LLaMA [59]. Because these LLMs are pre-trained on corpora containing millions to billions of heterogeneous observations, they potentially encode latent biological or behavioral information that is rarely captured within a single RCT cohort [9]. Thus, when LLM-generated predictions are integrated into established causal inference methods in a statistically principled manner, we hypothesize that these AI-generated predictions can reduce the variance of treatment effect estimates and boost the power to detect treatment effect heterogeneity.
Equally important, LLMs are particularly well-suited to extracting rich information from the unstructured data already being collected in modern RCTs. This is because LLMs have the ability to analyze text, images, and audio, allowing researchers to capture subtle individual characteristics, lived experiences, and contextual signals that structured variables often miss. Together, by importing external knowledge and unlocking information inside unstructured data in the trial, LLMs open up a path to substantially improve the power and precision of causal analyses in RCTs, ultimately enabling more individualized, equitable, and effective intervention strategies.
@@
Yet, not only developing rigorous and statistically valid causal inference methods for RCTs that integrate LLMs remains a major challenge, but current causal inference methods in RCTs rarely exploit the full spectrum of unstructured data, both limiting statistical power and the insights attainable from RCTs. On the one hand, naively plugging LLM-based predictions into causal estimators can induce bias and lead to problematic statistical inference, as LLMs lack guarantees on prediction accuracy and may rely on spurious correlations. Therefore, any causal inference method that incorporates LLM-based predictions must be explicitly robust to potential prediction errors. On the other hand, existing causal analyses of RCTs rely almost exclusively on structured variables, and when unstructured inputs are considered, investigators often tend to distill them into hand-engineered or processed features, which may discard nuanced information or introduce human biases [32]. Recent multimodal LLMs integrate these modalities into a shared latent space [3, 38, 41], making them suited to expose unstructured data and strengthen causal inference in RCTs.
@@ 1.2 Contribution
CALM introduces a novel statistical framework for valid inference on a broad class of causal parameters by leveraging LLM prediction in randomized experiments. Section 2 focuses on its applications to mean potential outcomes, and Section 3 extends the framework to average and heterogeneous average treatment effects. We demonstrate that, even when LLM predictions are biased, CALM remains both consistent and more efficient than augmented inverse propensity weighting (AIPW) estimators. This advantage stems from the heterogeneous calibration strategy introduced in Section 2 (Steps 3 and 3 in particular), which both correct bias in LLM-generated predictions and leverages their useful signal to sharpen causal estimation precision. In parallel, CALM automatically incorporates information from both structured and unstructured data by leveraging counterfactual outcome predictions generated by LLMs.
@@
This manuscript also contributes to the literature on causal inference using potentially biased auxiliary information [4, 6, 18]. Specifically, we demonstrate in Theorem 1 of Section 4 that our CALM estimator remains consistent even when the zero-shot predictions are biased for the mean potential outcomes. This robustness arises from a novel estimating equation that incorporates residualized LLM predictions, which are mean-zero by construction. Our procedure relies on four nuisance function estimates: the conditional expectation of the potential outcomes (as in classical augmented inverse probability weighting method [6, 50, 51]), the conditional mean of the zero-shot LLM predictions (for the residualization step), and the conditional variance and covariance of the true and predicted potential outcomes (used to construct the optimal weighting scheme). By employing sample splitting and Neyman orthogonalization, we achieve asymptotically normal distributional approximations and valid statistical inference under mild $L_{2}$-type conditions on these nuisance function estimates.
In addition to its methodological and theoretical advances, our work provides systematic empirical evidence on the practical utility of CALM. Using synthetic populations calibrated to the BRIGHTEN study with structured and unstructured data (Section 5.2), we demonstrate that CALM yields more accurate estimates of causal effects, achieving lower bias and reduced variance compared to benchmark methods (Section 5.3, Figure 3). We further demonstrate that CALM remains effective in both zero- and few-shot settings without domain-specific fine-tuning, adapts well to heterogeneous covariate strata, and avoids the under-coverage issues that arise when LLM predictions are naively incorporated as covariates in benchmark methods (Figure 3). In the BRIGHTEN case study in Section 6, CALM not only yields tighter confidence intervals than benchmark methods, but also detects subgroup treatment effects (e.g., among female Hispanic participants) that benchmark methods miss, indicating its potential to enhance the discovery of treatment heterogeneity. Finally, robustness checks across multiple LLMs and diverse prompt-engineering strategies in Section 6 confirm that CALM’s empirical performance is stable to design choices, thereby reducing practical barriers to applying LLMs in real-world experiment analysis.
@@ 1.3 Related literature
Our work relates to the literature on improving the efficiency of randomized experiments through covariate adjustment. Classical approaches such as regression adjustment and stratification are well known to reduce variance without compromising statistical validity [20, 37], but they are limited in their ability to utilize unstructured data and thus may not achieve full efficiency. More recently, double/debiased machine learning was introduced [13, 23, 27], which leverages flexible machine learning together with cross-fitting to estimate nuisance functions under weaker conditions while still reaching the efficiency bound. This framework has since been extended to other parameters, including the conditional average treatment effect [54] and the quantile treatment effect [30]. However, existing flexible machine learning tools, such as random forests and boosting, are primarily limited to structured data and are not directly applicable to unstructured data [14, 22, 45].
@@
Our work is also motivated by recent advances in LLMs, which have demonstrated remarkable predictive performance across diverse tasks. Early models focused on natural language tasks, such as text classification, question answering, and summarization, using models like BERT [16] and GPT-2 [46]. More recent models, such as GPT-3, PaLM, and LLaMA, exhibit strong in-context learning abilities, encompassing zero-shot and few-shot prediction, that make them adaptable to new tasks with reduced supervision [10, 17, 59]. In scientific and biomedical domains, domain-adapted models such as BioGPT and Galactica leverage specialized corpora to improve relevance and accuracy [39, 56]. At the same time, multimodal LLMs such as CLIP, Flamingo, BLIP-2, LLaVA, and GPT-4V extend these capabilities beyond text, enabling joint reasoning over text, images, and audio [2, 35, 38, 41, 47]. Beyond language and vision, LLMs have also been extended for reasoning over structured data and complex decision-making [26, 70]. However, most LLM-based prediction pipelines place little emphasis on uncertainty quantification or statistical guarantees, which are crucial for causal inference in high-stakes biomedical settings.
@@ 1.4 Problem setup: Causal parameters in randomized experiments
In this section, we introduce the notation used throughout the manuscript and then define the causal parameters of interest. Suppose we have a randomized experiment or a randomized controlled trial (RCT) consisting of $n$ subjects. Following the Neyman-Rubin causal model [52], let $T\in\{1,\ldots,k\}$ denote the randomized treatment assignment, $Y(t)$ the potential outcome under treatment $t\in\{1,\ldots,k\}$, and $Y$ the observed outcome.
For each subject in the experiment, we can observe pre-treatment covariate information. Let $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represent structured pre-treatment covariates that are routinely adopted for classical statistical analysis in randomized experiments, and let $Z\in\mathcal{Z}$ denote any unstructured pre-treatment covariates that are often challenging to incorporate. We denote the observed experimental data as $\{O_{i}\}_{i=1}^{n}$, where $O_{i}:=(Y_{i},T_{i},X_{i},Z_{i})$.
@@
This manuscript aims to provide valid and efficient statistical inference for several causal parameters, including (1) mean effect of treatment $t\in\{1,\ldots,k\}$:
$\mu_{t}=\mathbb{E}[Y(t)];$
(2) the average treatment effect (ATE) between treatment arms $t$ and $t^{\prime}$,
$\tau_{t,t^{\prime}}=\mathbb{E}[Y(t)-Y(t^{\prime})],\quad t\neq t^{\prime},\quad t,t^{\prime}\in\{1,\ldots,k\};$
(3) the conditional average treatment effect (CATE) given covariates $X=x$, for $x\in\mathcal{X}$:
$\tau_{t,t^{\prime}}(x)=\mathbb{E}[Y(t)-Y(t^{\prime})\mid X=x],$
@@
2 CALM for estimating $\mathbb{E}[Y(t)]$
In this section, we provide a detailed description along with simple heuristics of our proposed Causal Analysis leveraging Language Models (CALM) framework for $\mathbb{E}[Y(t)]$, leveraging either zero-shot or few-shot learning with pre-trained LLMs. In section 3, we extend CALM to ATE and CATE estimation.
@@ 2.1 CALM with zero-shot learning
We start with describing the step-by-step procedure for estimating $\mathbb{E}[Y(t)]$ using CALM with zero-shot learning. Alongside the detailed steps, we also provide simple heuristics and implementation guidance for practical use.
@@
More generally, one can also employ $K$-fold sample splitting; we use $K=2$ here for illustration. As we will demonstrate later (Section 4.1), this sample-splitting procedure relaxes the need for the Donsker condition when estimating conditional moments, thereby allowing the use of more flexible estimation methods [13].
@@
To provide details for Step 3, we clarify the definition of $\mu_{t}^{\dagger}(x)$, offer heuristic interpretations of $\omega_{t}(x)$, and outline their respective estimation strategies. To start with, $Y^{\dagger}(t)$ can be viewed as a function of structured covariates $X$ and unstructured covariates $Z$, denoted by $f_{\theta,t}(X,Z)$. Here, $\theta$ contains the pre-trained parameters of the LLM used in Step 1, and $f_{\theta,t}(\cdot)$ represents the mapping implemented by the LLM. The conditional expectation $\mu_{t}^{\dagger}(x):=\mathbb{E}\left(Y^{\dagger}(t)\mid X=x\right)$ is therefore taken with respect to $Z$, conditional on $X$. A more detailed discussion is provided in Section 4.
@@
Step 4 (CALM estimator for $\mathbb{E}[Y(t)]$)

### v2

@@ Abstract
Randomized experiments or randomized controlled trials (RCTs) are gold standards for causal inference, yet cost and sample-size constraints limit power. We introduce CALM (Causal Analysis leveraging Language Models), a statistical framework that integrates large language models (LLMs) generated insights of RCTs with established causal estimators to increase precision while preserving statistical validity. In particular, CALM treats LLM-generated outputs as auxiliary prognostic information and corrects their potential bias via a heterogeneous calibration step that residualizes and optimally reweights predictions. We prove that CALM remains consistent even when LLM predictions are biased and achieves efficiency gains over augmented inverse probability weighting estimators for various causal effects. In particular, CALM develops a few-shot variant that aggregates predictions across randomly sampled demonstration sets. The resulting U-statistic-like predictor restores i.i.d. structure and also mitigates prompt-selection variability. Empirically, in simulations calibrated to a mobile-app depression RCT, CALM delivers lower variance relative to other benchmarking methods, is effective in zero- and few-shot settings, and remains stable across prompt designs. By principled use of LLMs to harness unstructured data and external knowledge learned during pretraining, CALM provides a practical path to more precise causal analyses in RCTs.
Keywords: Artificial intelligence; Clinical trials; Few-shot learning.
@@ 1.1 Background and motivation
Reliably designed randomized controlled trials (RCTs) are widely regarded as the gold standard for evaluating the effectiveness of biomedical, behavioral, socioeconomic or business interventions. By leveraging randomization and controlled conditions, RCTs minimize bias and facilitate the credible estimation of causal effects. Additionally, subgroup analyses are frequently employed to explore treatment effect heterogeneity, thereby supporting the advancement of individualized treatment targeting. However, most datasets from RCTs are analyzed in isolation, without incorporating external information. Moreover, standard estimation strategies for average and subgroup effects often rely on regression analyses using only a few hand-picked pre-intervention covariates. This approach leaves significant baseline information unexploited, particularly variables of multi-modality nature or those subject to substantial missingness. These limitations collectively compromise statistical efficiency, a critical drawback given that RCTs are often constrained by their sample sizes due to high implementation costs.
@@
Recent advances in artificial intelligence (AI), large-scale pre-trained large language models (LLMs) in particular, offer potential opportunities to boost the statistical power of RCTs by leveraging the external knowledge obtained during training. LLMs are deep neural networks trained with self-supervision on broad and heterogeneous corpora, producing general representations that can be rapidly adapted to many downstream tasks with little or no additional labeled data [6]. Importantly, multimodal LLMs now span the biomedical and behavioral spectrum, with representative examples including GPT-4 [39], Gemini [21], and LLaMA [51], meaning that they potentially encode latent biological or behavioral information about the treatments [6, 22, 34]. When LLM-generated predictions are integrated into established causal inference methods in a statistically principled manner, we hypothesize that this can reduce the variance of treatment effect estimates and boost the power to detect treatment effect heterogeneity.
As an added benefit, LLMs are naturally well-suited for extracting rich information from the unstructured data increasingly collected in modern RCTs. Unstructured data, such as clinical notes, medical images, participant narratives, and transcripts of subject-provider interactions, are now standard in RCTs across oncology, mental health [42, 46], dementia [23], and healthcare delivery [8, 24] (see Table 1 in the Supplementary Materials for motivating examples). Historically, incorporating such data into causal analyses has often required ad hoc feature-engineering pipelines that can discard nuanced information, introduce human biases [27], or create additional high-dimensional covariates. By contrast, multimodal LLMs offer a more flexible and principled way to process these complex unstructured inputs.
@@
Yet, developing rigorous and statistically valid causal inference methods for RCTs that integrate LLM outputs remains a major challenge. This is because naively plugging LLM-based predictions into causal estimators can induce bias and lead to problematic statistical inference, as LLMs lack guarantees on prediction accuracy and may rely on spurious correlations. Therefore, any causal inference method that incorporates LLM-based predictions must be explicitly robust to potential prediction errors.
@@ 1.2 Contribution
CALM introduces a novel statistical framework for valid inference on a broad class of causal parameters by leveraging LLM prediction in randomized experiments. Section 2 focuses on its applications to mean potential outcomes, and Section 3 extends the framework to average and heterogeneous average treatment effects. We demonstrate that, even when LLM predictions are biased, CALM remains both consistent and more efficient than augmented inverse propensity weighting (AIPW) estimators. This advantage stems from the heterogeneous calibration strategy introduced in Section 2 (Steps 3 and 3 in particular), which both corrects bias in LLM-generated predictions and leverages LLM’s intrinsic knowledge and serialized input data to sharpen causal estimation precision.
@@
This manuscript also contributes to the literature on causal inference using potentially biased auxiliary information [4, 16, 35, 57]. Specifically, we demonstrate in Theorem 1 of Section 4 that our CALM estimator remains consistent even when the zero-shot predictions are biased for the mean potential outcomes. This robustness arises from a novel estimating equation that incorporates residualized LLM predictions, which are mean-zero by construction. Our procedure relies on four nuisance function estimates: the conditional expectation of the potential outcomes (as in classical augmented inverse probability weighting method [9, 17, 37]), the conditional mean of the zero-shot LLM predictions (for the residualization step), and the conditional variance and covariance of the true and predicted potential outcomes (used to construct the optimal weighting scheme). By employing sample splitting and Neyman orthogonalization, we achieve asymptotically normal distributional approximations and valid statistical inference under mild $L_{2}$-type conditions on these nuisance function estimates.
In addition to its methodological and theoretical advances, our work provides systematic empirical evidence on the practical utility of CALM. Using synthetic populations calibrated to the BRIGHTEN study with structured and unstructured data (Section 5.2), we demonstrate that CALM yields more accurate estimates of causal effects, achieving lower bias and reduced variance compared to benchmark methods (Section 5.3, Figure 3). In particular, CALM’s efficiency gain is independent of whether unstructured data are used. This indicates that the improvement does not stem from incorporating additional prognostic information from unstructured inputs, but rather from the intrinsic information gain provided by the LLM itself (Section 5.4, Figure 4). We further demonstrate that CALM remains effective in both zero- and few-shot settings without domain-specific fine-tuning, adapts well to heterogeneous covariate strata (Section 5.4, Figure 5), and avoids the under-coverage issues that arise when LLM predictions are naively incorporated as covariates in benchmark methods (Figure 3). In the BRIGHTEN case study in Section 6, CALM not only yields tighter confidence intervals than benchmark methods, but also detects subgroup treatment effects (e.g., among female Hispanic participants) that benchmark methods miss, indicating its potential to enhance the discovery of treatment heterogeneity. Finally, robustness checks across multiple LLMs and diverse prompt-engineering strategies in Section 6 confirm that CALM’s empirical performance is stable to design choices, thereby reducing practical barriers to applying LLMs in real-world experiment analysis.
@@ 1.3 Related literature
Our work relates to the literature on improving the efficiency of randomized experiments through covariate adjustment. Classical approaches such as regression adjustment and stratification are well known to reduce variance without compromising statistical validity [19, 32], but they are limited in their ability to utilize unstructured data and thus may not achieve full efficiency. More recently, double/debiased machine learning [11] leverages flexible machine learning together with cross-fitting. However, existing flexible machine learning tools, such as random forests and boosting, are primarily limited to structured data and are not directly applicable to unstructured data [12, 23, 42]. In addition, the inclusion of many covariates may lead to biases in the asymptotic normal distribution and therefore renders standard normal-based inference invalid [10].
@@
Our work is also motivated by recent advances in LLMs, which have demonstrated remarkable predictive performance across diverse tasks. Early models focused on natural language tasks, such as text classification, question answering, and summarization, using models like BERT [14] and GPT-2 [44]. More recent models, such as GPT-3, PaLM, and LLaMA, exhibit strong in-context learning abilities, encompassing zero-shot and few-shot prediction, that make them adaptable to new tasks with reduced supervision [7, 15, 51]. In scientific and biomedical domains, domain-adapted models such as BioGPT and Galactica leverage specialized corpora to improve relevance and accuracy [36, 49]. At the same time, multimodal LLMs such as CLIP, Flamingo, BLIP-2, LLaVA, and GPT-4V extend these capabilities beyond text, enabling joint reasoning over text, images, and audio [3, 30, 33, 39, 43]. However, most LLM-based prediction pipelines place little emphasis on uncertainty quantification or statistical guarantees, which are crucial for causal inference in high-stakes biomedical settings.
@@ 1.4 Problem setup: Causal parameters in randomized experiments
In this section, we introduce the notation used throughout the manuscript and then define the causal parameters of interest. Suppose we have a randomized experiment or a randomized controlled trial (RCT) consisting of $n$ subjects. Following the Neyman-Rubin causal model, let $T\in\{1,\ldots,k\}$ denote the randomized treatment assignment, $Y(t)$ the potential outcome under treatment $t\in\{1,\ldots,k\}$, and $Y$ the observed outcome.
For each subject in the experiment, we can observe pre-treatment covariate information. Let $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represent structured pre-treatment covariates that are routinely adopted for classical statistical analysis in randomized experiments, and let $Z\in\mathcal{Z}$ denote any variables that may contain heavy missing or unstructured pre-treatment data that are traditionally challenging to incorporate. We note that CALM’s efficiency gain does not depend on the presence of $Z$ and its ability to conveniently handle unstructured data is simply an additional benefit. We denote the observed experimental data as $\{O_{i}\}_{i=1}^{n}$, where $O_{i}:=(Y_{i},T_{i},X_{i},Z_{i})$.
@@
This manuscript aims to provide valid and efficient statistical inference for several causal parameters, including (i) mean effect of treatment $t\in\{1,\ldots,k\}$:
$\mu_{t}=\mathbb{E}(Y(t));$
(ii) the average treatment effect (ATE) between treatment arms $t$ and $t^{\prime}$,
$\tau_{t,t^{\prime}}=\mathbb{E}(Y(t)-Y(t^{\prime})),\quad t\neq t^{\prime},\quad t,t^{\prime}\in\{1,\ldots,k\};$
(iii) the conditional average treatment effect (CATE) given covariates $X=x$, for $x\in\mathcal{X}$:
$\tau_{t,t^{\prime}}(x)=\mathbb{E}(Y(t)-Y(t^{\prime})\mid X=x),$
@@
2 CALM for estimating $\mathbb{E}(Y(t))$
In this section, we provide a detailed description along with simple heuristics of our proposed Causal Analysis leveraging Language Models (CALM) framework for $\mathbb{E}(Y(t))$, leveraging either zero-shot or few-shot learning with pre-trained LLMs. In section 3, we extend CALM to ATE and CATE estimation.
@@ 2.1 CALM with zero-shot learning
We start with describing the step-by-step procedure for estimating $\mathbb{E}(Y(t))$ using CALM with zero-shot learning. Alongside the detailed steps, we also provide simple heuristics and implementation guidance for practical use.
@@
More generally, one can also employ $K$-fold sample splitting; we use $K=2$ here for illustration.
@@
To provide details for Step 3, we clarify the definition of $\mu_{t}^{\dagger}(x)$, offer heuristic interpretations of $\omega_{t}(x)$, and outline their respective estimation strategies. To start with, $Y^{\dagger}(t)$ can be viewed as a function of structured covariates $X$ and unstructured covariates $Z$, denoted by $f_{\theta,t}(X,Z)$. Here, $\theta$ contains the pre-trained parameters of the LLM used in Step 1, and $f_{\theta,t}(\cdot)$ represents the mapping implemented by the LLM. The conditional expectation $\mu_{t}^{\dagger}(x):=\mathbb{E}(Y^{\dagger}(t)\mid X=x)$ is therefore taken with respect to $Z$, conditional on $X$. A more detailed discussion is provided in Section 4.
@@
Step 4 (CALM estimator for $\mathbb{E}(Y(t))$)